A body moving with uniform acceleration, after successive intervals of time, is at A, B, C, D which are collinear and at distances a, b, c, d from some arbitrary point in the same straight line. Initially body is at rest at the arbitrary point. Then, (da) is equal to:

Grade:11

1 Answers

Agrata Singh
5 years ago
Let the arbitrary point be O and the constant accn be p 
From second eqn of motion: S=ut+1/2pt2
And u=0 (body starts its motion at O)
OA=0+1/2pt2=a
OB=0+1/2p(2t)2=4a=b
OC=0+1/2p(3t)2=9a=c
OD=0+1/2p(4t)2=16a=d
 
Hence d-a= 16a-a = 15a